2y^2+4xy=x^2
Solve for y.
I cannot figure out what to even do here
what could the answer be?

Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

2y^2+4xy=x^2
Solve for y.
-------------
2y^2+4xy-x^2 = 0
y = 2x + sqrt(4x^2 + 8x^2)/2
y = 2x + x*sqrt(3)
y = x*(2 + sqrt(3))
y = x*(2 - sqrt(3))

Answer by jim_thompson5910(35256)